OpenStudy (anonymous):

Find the equation of the line containing the given pair of components (-4,-7) and (-2,-1)

OpenStudy (luigi0210):

Find the slope first \(\large {m=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}}\) then use the point-slope formula \(\large y-y_{1}=m(x-x_{1})\)